# Homotopy connectivity of Čech complexes of spheres

### v1

@@ 5. An upper bound on connectivity
$\sigma\in N(\mathrm{Bor}(S^{n};\delta))\Longleftrightarrow$ $\exists\hskip 2.84526ptv\in S^{n}\text{ with }x\sim v\text{ in }\mathrm{Bor}(S^{n};\delta)\text{ for all }x\in\sigma$
@@
Let us illustrate Lovász’s result for the Petersen graph. The Petersen graph ${\mathcal{P}}$ is the Kneser graph $KG_{5,2}$, whose vertices are $2$-element subsets of a $5$-element set and two of them are joined by an edge if and only if they are disjoint. The degree of each vertex of ${\mathcal{P}}$ is $3$, and therefore the dimension of the the neighborhood complex $N({\mathcal{P}})$ is $2$. Each vertex of ${\mathcal{P}}$ gives rise to a unique triangle in $N({\mathcal{P}})$, and no two triangles in $N({\mathcal{P}})$ share an edge. Also, each vertex of $N({\mathcal{P}})$ is contained in exactly $3$ such triangles in $N({\mathcal{P}})$. It turns out that $N({\mathcal{P}})$ is homotopy equivalent to a finite connected graph with holes, meaning that $N({\mathcal{P}})$ is not simply connected and $\operatorname{conn}(N({\mathcal{P}}))=0$. Lovász’s bound [30, Theorem 2] gives $\chi({\mathcal{P}})\geq\operatorname{conn}(N({\mathcal{P}}))+3=0+3$, and in fact, $\chi({\mathcal{P}})=3$ by [30, Theorem 1’].
@@
The problem of bounding $\chi(\mathrm{Bor}(S^{n};\delta))$ for $0<\delta<\pi$ close to $\pi$ is classical. One version of the Borsuk–Ulam theorem (the Lyusternik–Schnirel’man–Borsuk covering theorem) gives, and is in fact equivalent to, the statement that $\chi(\mathrm{Bor}[S^{n};\delta])\geq n+2$ for all $\delta<\pi$; see, for example, [12, 29, 6] and [31, Section 2.1, Exercise 10]. Taking $\varepsilon>0$ such that $\delta+\varepsilon<\pi$, the left inclusion in (1) gives that $\chi(\mathrm{Bor}(S^{n};\delta))\geq n+2$, as well. Furthermore, Raigorodskii shows that $\chi(\mathrm{Bor}[S^{n};\delta])=n+2$ for all $s_{n}<\delta<\pi$, where $s_{n}$ is the diameter of the radial projection of a single $n$-dimensional face of a regular $(n+1)$-dimensional simplex inscribed in $S^{n}$ [37, 38]. Once again, (1) implies that $\chi(\mathrm{Bor}(S^{n};\delta))=n+2$ for all $s_{n}<\delta<\pi$; see Figure 3 for the case $n=2$. From [39] (after plugging $\cos\ell=-\frac{1}{n+1}$ into (2.17) and (2.18) of [39]), we have that
@@
The asymptotics of the chromatic numbers of Borsuk graphs of $S^{n}$ (though not the exact values) may be reasonably well controlled by covering and packing numbers on spheres. See, for example, [36, Theorem 5.1.5]222 Moy defines his covering radius $c_{n,m}$ using balls of radius $\frac{r}{2}$, whereas we define our covering radius $\operatorname{cov}_{S^{n}}(m)$ using balls of radius $r$, giving the relationship $c_{n,m}=2\operatorname{cov}_{S^{n}}(m)$., which states that when $\delta>2\operatorname{cov}_{S^{n}}(m)$ for some $m\geq 2$, then $\chi(\mathrm{Bor}[S^{n};\delta])\leq m$. Since $\mathrm{Bor}(S^{n};\delta)\subseteq\mathrm{Bor}[S^{n};\delta]$, we have that $\chi(\mathrm{Bor}(S^{n};\delta))\leq m$.
@@
Lovász’s upper bound in Lemma 5.2 gives $\operatorname{conn}(\mathrm{\check{C}}(S^{n};\pi-\delta))\leq\chi(\mathrm{Bor}(S^{n};\delta))-3$. When $\delta>2\operatorname{cov}_{S^{n}}(k+1)$ for some $k+1\geq 2$, then $\chi(\mathrm{Bor}(S^{n};\delta))\leq k+1$ by [36, Theorem 5.1.5]. Together, these give
@@ 6. The case of the circle
In the case of the circle, we know the homotopy types of $\mathrm{\check{C}}(S^{1};r)$ for all $r$; see [1, Section 9]. This gives us one “data point” ($n=1$) that any more general results for $\mathrm{\check{C}}(S^{n};r)$ with $n\geq 1$ will need to generalize. See Figure 2.
By [1, Theorem 9.8], the homotopy types of $\mathrm{\check{C}}(S^{1};r)$ are all odd spheres. It turns out that
@@ 6.1. Comparing the lower bound
From [1, Theorem 7.4], we have that
[…]
@@ 6.2. Comparing the upper bound
$\chi(\mathrm{Bor}(S^{1};\delta))=\lceil\frac{2\pi}{\delta}\rceil$ for any $0<\delta<\pi$.
@@
We know that $\chi(\mathrm{Bor}[S^{1};\delta])=\lceil\frac{2\pi}{\delta}\rceil$; see, for example, [36, Section 5.1.2]. So, for $k\geq 2$, $\chi(\mathrm{Bor}[S^{1};\delta])=k$ when $\tfrac{2\pi}{k}\leq\delta<\tfrac{2\pi}{k-1}$. Let $\tfrac{2\pi}{k}\leq\delta<\tfrac{2\pi}{k-1}$, and choose $\varepsilon>0$ such that $\delta+\varepsilon<\tfrac{2\pi}{k-1}$. Then by (1), $k=\chi(\mathrm{Bor}[S^{1};\delta+\varepsilon])\leq\chi(\mathrm{Bor}(S^{1};\delta))\leq\chi(\mathrm{Bor}[S^{1};\delta])=k$. Thus, we have that $\chi(\mathrm{Bor}(S^{1};\delta))=\chi(\mathrm{Bor}[S^{1};\delta])=\lceil\frac{2\pi}{\delta}\rceil$. ∎
@@
  • Plugging $n=1$ into Lovász’s bound $\operatorname{conn}(\mathrm{\check{C}}(S^{n};\pi-\delta))\leq\chi(\mathrm{Bor}(S^{n};\delta))-3$, and using the exact chromatic numbers of the Borsuk graphs $\mathrm{Bor}(S^{1};\delta)$ from Lemma 6.2, together give $\operatorname{conn}(\mathrm{\check{C}}(S^{1};\pi-\delta))\leq\lceil\frac{2\pi}{\delta}\rceil-3$. If $\operatorname{conn}(\mathrm{\check{C}}(S^{1};\pi-\delta))=2k$, then $2k+3\leq\lceil\frac{2\pi}{\delta}\rceil$, which implies that $\delta<\frac{2\pi}{2k+2}=\frac{\pi}{k+1}$.
  • Proposition 5.3 states if $\operatorname{conn}(\mathrm{\check{C}}(S^{n};\pi-\delta))\geq k-1$, then $\delta\leq 2\operatorname{cov}_{S^{n}}(k+1)$. For $n=1$, if $\operatorname{conn}(\mathrm{\check{C}}(S^{1};\pi-\delta))\geq k-1$, then $\delta\leq 2\operatorname{cov}_{S^{1}}(k+1)=\frac{2\pi}{k+1}$. Via the substitution $k\mapsto 2k+1$, we get that if $\operatorname{conn}(\mathrm{\check{C}}(S^{1};\pi-\delta))\geq 2k$, then $\delta\leq\frac{2\pi}{2k+2}=\frac{\pi}{k+1}$.
  • By (3), $\operatorname{conn}(\mathrm{\check{C}}(S^{1};\pi-\delta)=2k$ if and only if $\delta\in[\frac{\pi}{k+2},\frac{\pi}{k+1})$.
Lovász’s bound $\operatorname{conn}(\mathrm{\check{C}}(S^{n};\pi-\delta))\leq\chi(\mathrm{Bor}(S^{n};\delta))-3$ and Proposition 5.3 both give the optimal upper bound of $\frac{\pi}{k+1}$ on $\delta$, for even connectivity, when $n=1$. The former excludes the possibility $\delta=\frac{\pi}{k+1}$, whereas the latter does not, and hence, the former upper bound is slightly better. This makes sense since our Proposition 5.3 is derived from Lovász’s bound but does not use exact chromatic numbers of Borsuk graphs.
@@ 7. Homological dimension
The ${\mathbb{Z}}$-homological dimension of a space is informative of its topology: a space that is homotopy equivalent to a CW complex of dimension $k$ has ${\mathbb{Z}}$-homological dimension at most $k$. The ${\mathbb{Z}}$-homological dimension of a space is often connected to its homotopy connectivity. Indeed, if $Y$ is simply connected, then by the Hurewicz theorem [22, Theorem 4.32], we have $\operatorname{conn}(Y)\leq\operatorname{h-dim}_{{\mathbb{Z}}}(Y)-1$. Note that $\operatorname{conn}(Y)$ need not333 To see that we may have $\operatorname{h-dim}_{{\mathbb{Z}}}(X)>\operatorname{conn}(X)>\operatorname{h-dim}_{{\mathbb{Z}}_{2}}$, let $p$ be an odd prime, and let $X=M({\mathbb{Z}}_{p},r+1)$ be a Moore space that is $r$-connected for some $r>0$. Then $\operatorname{h-dim}_{{\mathbb{Z}}}(X)=r+1>r>0=\operatorname{h-dim}_{{\mathbb{Z}}_{2}}(X)$. be a lower bound to $\operatorname{h-dim}_{R}(Y)-1$ for a general ring $R$.
We know that $\mathrm{\check{C}}(S^{n};\pi-\delta)\simeq S^{n}$ for all $\delta\in[\pi/2,\pi)$ due to the nerve lemma. We also know that $\mathrm{\check{C}}(S^{1};\pi-\delta)$ is homotopic to some odd-dimensional sphere for each $\delta>0$. Hence, the ${\mathbb{Z}}$-homological dimension of $\mathrm{\check{C}}(S^{n};\pi-\delta)$ is known in all these cases. So, let us restrict our attention to the case when $n\geq 2$ and $\delta\in(0,\pi/2)$.
In Proposition 4.1, we gave a lower bound on the connectivity $\operatorname{conn}(\mathrm{\check{C}}(S^{n};\pi-\delta))$. In particular, if $0<\delta<\operatorname{cov}_{S^{n}}(2k+2)$, then $\operatorname{conn}(\mathrm{\check{C}}(S^{n};\pi-\delta))\geq k$. For $n\geq 2$, it follows from [45, Theorem 11.2] that the Čech complex $\mathrm{\check{C}}(S^{n};\pi-\delta)$ is simply connected for each $\delta\in(0,\pi/2)$. Hence, using the relation between connectivity and ${\mathbb{Z}}$-homological dimension stated in the previous paragraph, we get a lower bound to $\operatorname{h-dim}_{{\mathbb{Z}}}(\mathrm{\check{C}}(S^{n};\pi-\delta))$: if $0<\delta<\operatorname{cov}_{S^{n}}(2k+2)$, then $\operatorname{h-dim}_{{\mathbb{Z}}}(\mathrm{\check{C}}(S^{n};\pi-\delta))\geq k+1$. In other words, for any $\delta\in(0,\pi/2)$ and $n\geq 2$, Corollary 4.3 gives that
@@
There are several related definitions of box complexes, which all share similar properties. For a finite graph $G$, the box complex ${\mathcal{B}}(G)$ is a free ${\mathbb{Z}}_{2}$-simplicial complex satisfying the following properties:
@@
To bound $\operatorname{h-dim}_{{\mathbb{Z}}_{2}}(\mathrm{\check{C}}(X,S^{n};r))$ from below for finite $X\subseteq S^{n}$, we will also use the fact that if $Y$ is a finite free ${\mathbb{Z}}_{2}$-simplicial complex and if there is a ${\mathbb{Z}}_{2}$-map from $S^{k}$ to $Y$, then $\operatorname{h-dim}_{{\mathbb{Z}}_{2}}(Y)\geq k$; see [32, Lemma 7] or [33, Lemma 4.4]. We also use the packing number from Section 3.1.
@@
It seems that the proof technique of [32, Lemma 7] might offer tools towards proving (5) (and consequently (6)). Indeed, proceeding as in the proof of [32, Lemma 7], we get that for an infinite ${\mathbb{Z}}_{2}$-complex $Y$ equipped with a ${\mathbb{Z}}_{2}$-map $S^{k}\to Y$, either $\operatorname{h-dim}_{{\mathbb{Z}}_{2}}(Y)\geq k$ or $\operatorname{h-dim}_{{\mathbb{Z}}_{2}}(\tfrac{Y}{{\mathbb{Z}}_{2}})$ is infinite.
@@
Let us see how good of a bound (6) would be, in the case of the Čech complexes of a circle. By  [1, Theorem 9.8] and (2), we have
@@
Also, note that $\operatorname{numPack}_{S^{1}}(\delta)=k$ when $\delta\in[\tfrac{2\pi}{k+1},\tfrac{2\pi}{k})$. Thus, if $\operatorname{h-dim}_{{\mathbb{Z}}}(\mathrm{\check{C}}(S^{1};\pi-\delta))\leq 2k+1$, then according to (6), we would have $\operatorname{numPack}_{S^{n}}(\delta)\leq 2k+3$. This implies that $\delta\geq\frac{2\pi}{2k+4}$ or $\delta\geq\frac{\pi}{k+2}$, which is exactly the lower bound given by  [1, Theorem 9.8]. Hence, the lower bound in (6) would be tight in the case of the circle, and also a $4$-fold improvement over the prior lower bound (4). However, we do not know whether results similar to that of [33, Lemma 4.3] hold for infinite graphs.

### v2

@@ 5. An upper bound on connectivity
$\sigma\in N(\mathrm{Bor}(S^{n};\delta))\Longleftrightarrow\$$\exists\hskip 2.84526ptv\in S^{n}\text{ with }x\sim v\text{ in }\mathrm{Bor}(S^{n};\delta)\text{ for all }x\in\sigma$
@@
Let us illustrate Lovász’s result for the Petersen graph. The Petersen graph $\mathcal{P}$} is the Kneser graph $KG_{5,2}$, whose vertices are $2$-element subsets of a $5$-element set and two of them are joined by an edge if and only if they are disjoint. The degree of each vertex of $\mathcal{P}$} is $3$, and therefore the dimension of the neighborhood complex $N({\mathcal{P}})$ is $2$. Each vertex of $\mathcal{P}$} gives rise to a unique triangle in $N({\mathcal{P}})$, and no two triangles in $N({\mathcal{P}})$ share an edge. Also, each vertex of $N({\mathcal{P}})$ is contained in exactly $3$ such triangles in $N({\mathcal{P}})$. It turns out that $N({\mathcal{P}})$ is homotopy equivalent to a finite connected graph with holes, meaning that $N({\mathcal{P}})$ is not simply connected and $\operatorname{conn}(N({\mathcal{P}}))=0$. Lovász’s bound [30, Theorem 2] gives $\chi({\mathcal{P}})\geq\operatorname{conn}(N({\mathcal{P}}))+3=0+3$, and in fact, $\chi({\mathcal{P}})=3$ by [30, Theorem 1’].
@@
The problem of bounding $\chi(\mathrm{Bor}(S^{n};\delta))$ for $0<\delta<\pi$ close to $\pi$ is classical. One version of the Borsuk–Ulam theorem (the Lyusternik–Schnirel’man–Borsuk covering theorem) gives, and is in fact equivalent to, the statement that $\chi(\mathrm{Bor}[S^{n};\delta])\geq n+2$ for all $\delta<\pi$; see, for example, [12, 29, 6] and [32, Section 2.1, Exercise 10]. Taking $\varepsilon>0$ such that $\delta+\varepsilon<\pi$, the left inclusion in (1) gives that $\chi(\mathrm{Bor}(S^{n};\delta))\geq n+2$, as well. Furthermore, Raigorodskii shows that $\chi(\mathrm{Bor}[S^{n};\delta])=n+2$ for all $s_{n}<\delta<\pi$, where $s_{n}$ is the diameter of the radial projection of a single $n$-dimensional face of a regular $(n+1)$-dimensional simplex inscribed in $S^{n}$ [38, 39]. Once again, (1) implies that $\chi(\mathrm{Bor}(S^{n};\delta))=n+2$ for all $s_{n}<\delta<\pi$; see Figure 3 for the case $n=2$. From [40] (after plugging $\cos\ell=-\frac{1}{n+1}$ into (2.17) and (2.18) of [40]), we have that
@@
The asymptotics of the chromatic numbers of Borsuk graphs of $S^{n}$ (though not the exact values) may be reasonably well controlled by covering and packing numbers on spheres. See, for example, [37, Theorem 5.1.5] [^2] Moy defines his covering radius $c_{n,m}$ using balls of radius $\frac{r}{2}$, whereas we define our covering radius $\operatorname{cov}_{S^{n}}(m)$ using balls of radius $r$, giving the relationship $c_{n,m}=2\operatorname{cov}_{S^{n}}(m)$., which states that when $\delta>2\operatorname{cov}_{S^{n}}(m)$ for some $m\geq 2$, then $\chi(\mathrm{Bor}[S^{n};\delta])\leq m$. Since $\mathrm{Bor}(S^{n};\delta)\subseteq\mathrm{Bor}[S^{n};\delta]$, we have that $\chi(\mathrm{Bor}(S^{n};\delta))\leq m$.
@@
We recall that $\delta\in(0,\pi)$ in this paper. This excludes the possibility $k=1$ in Proposition 5.3, which means that the upper bound on $\operatorname{conn}(\mathrm{\check{C}}(S^{n};\pi-\delta))$ given above is always non-negative. Indeed, if $k=1$, then $2\operatorname{cov}_{S^{n}}(k+1)=\pi$ and we get the contradicting inequality $\pi-\delta<0$.
@@
Lovász’s upper bound in Lemma 5.2 gives $\operatorname{conn}(\mathrm{\check{C}}(S^{n};\pi-\delta))\leq\chi(\mathrm{Bor}(S^{n};\delta))-3$. When $\delta>2\operatorname{cov}_{S^{n}}(k+1)$ for some $k+1\geq 2$, then $\chi(\mathrm{Bor}(S^{n};\delta))\leq k+1$ by [37, Theorem 5.1.5]. Together, these give
@@ 6. The case of the circle
In the case of the circle, we know the homotopy types of $\mathrm{\check{C}}(S^{1};r)$ for all $r$; see [2, Section 9]. This gives us one “data point” ($n=1$) that any more general results for $\mathrm{\check{C}}(S^{n};r)$ with $n\geq 1$ will need to generalize. See Figure 2.
By [2, Theorem 9.8], the homotopy types of $\mathrm{\check{C}}(S^{1};r)$ are all odd spheres. It turns out that
@@ 6.1. Comparing the lower bound
From [2, Theorem 7.4], we have that
$\mathrm{VR}(S^{1};r)\simeq S^{2k+1}\hskip 14.22636pt\text{if}\hskip 14.22636ptr\in\left(\frac{2\pi k}{2k+1},\frac{2\pi(k+1)}{2k+3}\right].$
@@ 6.2. Comparing the upper bound
$\chi(\mathrm{Bor}(S^{1};\delta))=\bigl\lceil\frac{2\pi}{\delta}\bigr\rceil$ for any $0<\delta<\pi$.
@@
We know that $\chi(\mathrm{Bor}[S^{1};\delta])=\bigl\lceil\frac{2\pi}{\delta}\bigr\rceil$; see, for example, [37, Section 5.1.2]. So, for $k\geq 2$, $\chi(\mathrm{Bor}[S^{1};\delta])=k$ when $\tfrac{2\pi}{k}\leq\delta<\tfrac{2\pi}{k-1}$. Let $\tfrac{2\pi}{k}\leq\delta<\tfrac{2\pi}{k-1}$, and choose $\varepsilon>0$ such that $\delta+\varepsilon<\tfrac{2\pi}{k-1}$. Then by (1), $k=\chi(\mathrm{Bor}[S^{1};\delta+\varepsilon])\leq\chi(\mathrm{Bor}(S^{1};\delta))\leq\chi(\mathrm{Bor}[S^{1};\delta])=k$. Thus, we have that $\chi(\mathrm{Bor}(S^{1};\delta))=\chi(\mathrm{Bor}[S^{1};\delta])=\bigl\lceil\frac{2\pi}{\delta}\bigr\rceil$. ∎
@@
  • Plugging $n=1$ into Lovász’s bound $\operatorname{conn}(\mathrm{\check{C}}(S^{n};\pi-\delta))\leq\chi(\mathrm{Bor}(S^{n};\delta))-3$, and using the exact chromatic numbers of the Borsuk graphs $\mathrm{Bor}(S^{1};\delta)$ from Lemma 6.2, together give $\operatorname{conn}(\mathrm{\check{C}}(S^{1};\pi-\delta))\leq\bigl\lceil\tfrac{2\pi}{\delta}\bigr\rceil-3$. If $\operatorname{conn}(\mathrm{\check{C}}(S^{1};\pi-\delta))=2k$, then $2k+3\leq\bigl\lceil\tfrac{2\pi}{\delta}\bigr\rceil$, which implies that $\delta<\tfrac{2\pi}{2k+2}=\tfrac{\pi}{k+1}$.
  • Proposition 5.3 states if $\operatorname{conn}(\mathrm{\check{C}}(S^{n};\pi-\delta))\geq k-1$, then $\delta\leq 2\operatorname{cov}_{S^{n}}(k+1)$. For $n=1$, if $\operatorname{conn}(\mathrm{\check{C}}(S^{1};\pi-\delta))\geq k-1$, then $\delta\leq 2\operatorname{cov}_{S^{1}}(k+1)=\tfrac{2\pi}{k+1}$. Via the substitution $k\mapsto 2k+1$, we get that if $\operatorname{conn}(\mathrm{\check{C}}(S^{1};\pi-\delta))\geq 2k$, then $\delta\leq\tfrac{2\pi}{2k+2}=\tfrac{\pi}{k+1}$.
  • By (3), $\operatorname{conn}(\mathrm{\check{C}}(S^{1};\pi-\delta))=2k$ if and only if $\delta\in\bigl[\tfrac{\pi}{k+2},\tfrac{\pi}{k+1}\bigr)$.
Lovász’s bound $\operatorname{conn}(\mathrm{\check{C}}(S^{n};\pi-\delta))\leq\chi(\mathrm{Bor}(S^{n};\delta))-3$ and Proposition 5.3 both give the optimal upper bound of $\tfrac{\pi}{k+1}$ on $\delta$, for even connectivity, when $n=1$. The former excludes the possibility $\delta=\tfrac{\pi}{k+1}$, whereas the latter does not, and hence, the former upper bound is slightly better. This makes sense since our Proposition 5.3 is derived from Lovász’s bound but does not use exact chromatic numbers of Borsuk graphs.
@@ 7. Homological dimension
The $\mathbb{Z}$}-homological dimension of a space is informative of its topology: a space that is homotopy equivalent to a CW complex of dimension $k$ has $\mathbb{Z}$}-homological dimension at most $k$. The $\mathbb{Z}$}-homological dimension of a space is often connected to its homotopy connectivity. Indeed, if $Y$ is simply connected, then by the Hurewicz theorem [22, Theorem 4.32], we have $\operatorname{conn}(Y)\leq\operatorname{h-dim}_{{\mathbb{Z}}}(Y)-1$. Note that $\operatorname{conn}(Y)$ need not [^3] To see that we may have $\operatorname{h-dim}_{{\mathbb{Z}}}(X)>\operatorname{conn}(X)>\operatorname{h-dim}_{{\mathbb{Z}}_{2}}(X)$, let $p$ be an odd prime, and let $X=M({\mathbb{Z}}_{p},r+1)$ be a Moore space that is $r$-connected for some $r>0$. Then $\operatorname{h-dim}_{{\mathbb{Z}}}(X)=r+1>r>0=\operatorname{h-dim}_{{\mathbb{Z}}_{2}}(X)$. be a lower bound to $\operatorname{h-dim}_{R}(Y)-1$ for a general ring $R$.
We know that $\mathrm{\check{C}}(S^{n};\pi-\delta)\simeq S^{n}$ for all $\delta\in[\tfrac{\pi}{2},\pi)$ due to the nerve lemma. We also know that $\mathrm{\check{C}}(S^{1};\pi-\delta)$ is homotopic to some odd-dimensional sphere for each $\delta>0$. Hence, the $\mathbb{Z}$}-homological dimension of $\mathrm{\check{C}}(S^{n};\pi-\delta)$ is known in all these cases. So, let us restrict our attention to the case when $n\geq 2$ and $\delta\in(0,\tfrac{\pi}{2})$.
In Proposition 4.1, we gave a lower bound on the connectivity $\operatorname{conn}(\mathrm{\check{C}}(S^{n};\pi-\delta))$. In particular, if $0<\delta<\operatorname{cov}_{S^{n}}(2k+2)$, then $\operatorname{conn}(\mathrm{\check{C}}(S^{n};\pi-\delta))\geq k$. For $n\geq 2$, it follows from [46, Theorem 11.2] that the Čech complex $\mathrm{\check{C}}(S^{n};\pi-\delta)$ is simply connected for each $\delta\in(0,\tfrac{\pi}{2})$. Hence, using the relation between connectivity and $\mathbb{Z}$}-homological dimension stated in the previous paragraph, we get a lower bound to $\operatorname{h-dim}_{{\mathbb{Z}}}(\mathrm{\check{C}}(S^{n};\pi-\delta))$: if $0<\delta<\operatorname{cov}_{S^{n}}(2k+2)$, then $\operatorname{h-dim}_{{\mathbb{Z}}}(\mathrm{\check{C}}(S^{n};\pi-\delta))\geq k+1$. In other words, for any $\delta\in(0,\tfrac{\pi}{2})$ and $n\geq 2$, Corollary 4.3 gives that
@@
There are several related definitions of box complexes (see, for instance, [31, 17]), which all share similar properties. In particular, we are using the following definition of box complex. For a finite graph $G$, the box complex ${\mathcal{B}}(G)$ is defined to be
${\mathcal{B}}(G)=\{A\uplus B\,\mid\ A,B\subset V(G),\ A\cap B=\emptyset,\ G[A,B]\text{ is complete bipartite},\ \operatorname{CN}(A)\neq\emptyset\neq\operatorname{CN}(B)\}.$
Here, $A\uplus B=\{(a,0)\mid a\in A\}\cup\{(b,1)\mid b\in B\}$, $\operatorname{CN}(A)=\{v\in V(G)\mid(v,a)\in E(G)\ \text{for all }a\in A\}$, $\operatorname{CN}(B)=\{v\in V(G)\mid(v,b)\in E(G)\ \text{for all }b\in B\}$, and $G[A,B]$ is a (not necessarily induced) subgraph of $G$ with vertex set $A\cup B$ and edge set $\{(a,b)\in E(G)\mid a\in A,b\in B\}$, where $V(G)$ and $E(G)$ denote the sets of vertices and edges of $G$, respectively.
For a finite graph $G$, the box complex ${\mathcal{B}}(G)$ is a free ${\mathbb{Z}}_{2}$-simplicial complex satisfying the following three properties:
@@
To bound $\operatorname{h-dim}_{{\mathbb{Z}}_{2}}(\mathrm{\check{C}}(X,S^{n};r))$ from below for finite $X\subseteq S^{n}$, we will also use the fact that if $Y$ is a finite free ${\mathbb{Z}}_{2}$-simplicial complex and if there is a ${\mathbb{Z}}_{2}$-map from $S^{k}$ to $Y$, then $\operatorname{h-dim}_{{\mathbb{Z}}_{2}}(Y)\geq k$; see [34, Lemma 7] or [33, Lemma 4.4]. We also use the packing number from Section 3.1.
@@
It seems that the proof technique of [34, Lemma 7] might offer tools towards proving (5) (and consequently (6)). Indeed, proceeding as in the proof of [34, Lemma 7], we get that for an infinite ${\mathbb{Z}}_{2}$-complex $Y$ equipped with a ${\mathbb{Z}}_{2}$-map $S^{k}\to Y$, either $\operatorname{h-dim}_{{\mathbb{Z}}_{2}}(Y)\geq k$ or $\operatorname{h-dim}_{{\mathbb{Z}}_{2}}(\tfrac{Y}{{\mathbb{Z}}_{2}})$ is infinite.
@@
Let us see how good of a bound (6) would be, in the case of the Čech complexes of a circle. By  [2, Theorem 9.8] and (2), we have
@@
Also, note that $\operatorname{numPack}_{S^{1}}(\delta)=k$ when $\delta\in[\tfrac{2\pi}{k+1},\tfrac{2\pi}{k})$. Thus, if $\operatorname{h-dim}_{{\mathbb{Z}}}(\mathrm{\check{C}}(S^{1};\pi-\delta))\leq 2k+1$, then according to (6), we would have $\operatorname{numPack}_{S^{n}}(\delta)\leq 2k+3$. This implies that $\delta\geq\frac{2\pi}{2k+4}$ or $\delta\geq\frac{\pi}{k+2}$, which is exactly the lower bound given by  [2, Theorem 9.8]. Hence, the lower bound in (6) would be tight in the case of the circle, and also a $4$-fold improvement over the prior lower bound (4). However, we do not know whether results similar to that of [33, Lemma 4.3] hold for infinite graphs.